how do i rewrite the following using a single exponent?
4^(10m-6)/4^(9m-7)=4

Answer by stanbon(75887)

how do i rewrite the following using a single exponent?
4^(10m-6)/4^(9m-7)=4
-----
4^[10m-6-9m+7] = 4
=======
4^[m+1] = 4
----
= 4^m
===========
Cheers,
Stan H.